Suppose that you are a manager for MegaAccounting corporation, a firm specializing in accounting software. You know that you have two types of clients who use your software. Type A's inverse demand function is given by p=40-6q and type B's inverse demand function is given by p=37-11q. Your firm faces a constant marginal cost curve at $20

Suppose you can prevent buyers from trading with each other. What would be the price you would set in market A
Answer #1

Type A ( Market A)

p = 40 - 6q

TR = pq

= (40 - 6q)q

= 40q - 6q2

dTR/dq = MR = 40 - 12q

MR = 40 - 12q

Since buyer are preventing from trading with each other thus, manager will charge price according to MR = MC

40 - 12q = 20

40 - 20 = 12q

20 = 12q

q = 20/12

= 5/3

p = 40 - 6(5/3)

= 40 - 10  

= 30

thus, price that will be charged in market A is 30